OpenStudy (anonymous):

Write the solutions of each inequality in set-builder notation. 2(3p - 11) *greater than or equal to* -16

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

2(3p - 11) >= -16 --- distribute through the parenthesis 6p - 22 >= -16 --- add 22 to both sides 6p >= -16 + 22 6p >= 7 -- divide both sides by 6 p >= 7/6 or 1 1/6 { p | 7/6 < = p ......I am not exactly sure how to write this in set builder notation. p is gonna be any number greater or equal to 7/6

OpenStudy (texaschic101):

I think it might just be : { p | p >= 7/6 }

OpenStudy (anonymous):

YEAH !! Ur da best :)